### v1

# Analytical approach for calculating shadow of dynamical black hole

Abstract
A real astrophysical black hole, surrounded by an accretion disc, experiences time-dependent mass variation due to accretion, impacting observable features like the shadow. In this work, we analyzed the dynamical shadow by extending the concept of the photon sphere to time-dependent spacetimes, incorporating induced, centrifugal, and relativistic forces. We demonstrated that the shadow grows when light loses energy and shrinks when it gains energy, with a surprising result that the shadow remains static when the photon sphere radius coincides with the NEC horizon radius. Our method, assuming a small accretion rate, provides a foundation for studying dynamical black holes, offering insights into processes like gravitational collapse, which could make shadow evolution detectable and reveal details about black hole formation.
Black hole; Dynamical; Vaidya spacetime; Shadow; Photon sphere.
pacs:
95.30.Sf, 04.70.-s, 97.60.Lf, 04.50.Kd
I Introduction
In the early 20th century, general relativity proposed a range of phenomena, many of which were later validated through experimental observations. Among these were the anomalous precession of Mercury’s perihelion and the deflection of light rays by the Sun’s gravitational field. The anomalous precession arises due to the influence of general relativistic corrections incorporated into the equations of motion. These corrections, which scale as $\sim 1/r^{4}$, become particularly significant in regions close to the gravitating source.Consequently, these effects reach their peak near black holes, where, intriguingly, precession can be both positive and negative, as demonstrated in recent studies Bambhaniya et al. (2021). For light, the curvature of its trajectory intensifies with the mass of the source, leading to more pronounced bending. In extreme cases, such as around black holes, light can bend so strongly that it follows a circular orbit around the black hole. These orbits constitute the photon sphere, a region that produces a striking observational feature: a dark spot in the sky termed the black hole shadow. This shadow is approximately $2.5$ times larger than the event horizon, offering a direct visual signature of the black hole’s immense gravitational influence.
The observation of the black hole shadow has played a pivotal role in establishing black holes as tangible astrophysical objects. In 2019, the Event Horizon Telescope Collaboration captured the first images of the shadow of the supermassive black hole $M87^{*}$ Akiyama et al. (2019). This milestone was followed three years later by the observation of the shadow of Sagittarius $A^{*}$, located at the center of the Milky Way galaxy Akiyama et al. (2022). These groundbreaking discoveries significantly heightened the scientific community’s interest in black holes.
The foundational methods for observing black hole photon sphere and shadows were introduced in  Virbhadra and Ellis (2000); Claudel et al. (2001); Virbhadra and Keeton (2008); Virbhadra (2022); Falcke et al. (2000), and since then, techniques for calculating these shadows have been continuously refined Tsupko and Bisnovatyi-Kogan (2020); Perlick et al. (2015); Tsupko (2021); Vertogradov and Övgün (2024a, b); Adler and Virbhadra (2022). For a comprehensive overview of this field, see the detailed review in Perlick and Tsupko (2022).
The shadow of a black hole has been proposed as a powerful tool for cosmological studies, serving as a “cosmological ruler” to measure and test various phenomena Tsupko et al. (2020); Vagnozzi et al. (2023); Chen et al. (2022); Allahyari et al. (2020). Through its observation, different black hole models can be rigorously tested Kuang and Övgün (2022); Kuang et al. (2022); Meng et al. (2022); Wang et al. (2023); Vagnozzi et al. (2023); Abdujabbarov et al. (2016a); Atamurotov et al. (2013); Abdujabbarov et al. (2015); Atamurotov et al. (2021); Gralla et al. (2019); Cunha and Herdeiro (2018); Afrin et al. (2021); Abdujabbarov et al. (2016b); Younsi et al. (2016); Konoplya and Zinhailo (2020); Konoplya (2019); Kumar Walia (2024); Rodríguez et al. (2024); Wang et al. (2017); Pantig and Övgün (2023, 2023); Pantig et al. (2022, 2023); Kumar and Ghosh (2020a, b); Okyay and Övgün (2022); Heidari et al. (2024); Lambiase et al. (2024a, b); Övgün et al. (2024); Karshiboev et al. (2024); Puliçe et al. (2023); Kumaran and Övgün (2023); Chakhchi et al. (2025, 2024); Belhaj et al. (2021), and intriguingly, even singularities—regions of spacetime where densities become infinite—can cast shadows Shaikh et al. (2019); Heydarzade and Vertogradov (2024).
Most existing methods for analyzing black hole shadows, however, are focused on static black holes. In reality, astrophysical black holes are dynamic systems, typically surrounded by an accretion disk and increasing in mass due to ongoing accretion processes. This dynamic environment introduces complexities that necessitate further refinement of theoretical models to account for the interplay between the shadow, the accretion disk, and the evolving mass of the black hole. There are only few examples in literature Tan (2023); Koga et al. (2022); Solanki and Perlick (2022); Mishra et al. (2019); Heydarzade and Vertogradov (2024); Vertogradov (2024a)
Black holes can also lose mass through radiation, further underscoring the need for their description using a dynamical metric. However, introducing a time-dependent framework significantly complicates the analysis, as it abandons the symmetry provided by time-translation invariance.
In the case of dynamical black holes, many effects observed in static black holes become challenging to describe. Phenomena such as the Penrose process, high-energy particle collisions, the black hole shadow, and the negative energy problem Vertogradov (2020) rely heavily on the existence of well-defined geodesic structures, which are facilitated by a conformal Killing vector.
The conformal Killing vector has proven instrumental in exploring several critical aspects of black hole physics. For instance, it has been used to analyze the charged Penrose effect Vertogradov (2023), investigate high-energy collisions during gravitational collapse Vertogradov (2024b), and study the behavior of the black hole shadow Solanki and Perlick (2022); Heydarzade and Vertogradov (2024). These applications highlight the necessity of symmetry considerations for the theoretical exploration of dynamic black holes and their associated phenomena.
For dynamical black holes with general spherical symmetry, only one conserved quantity exists: the angular momentum per unit mass, $L$. In this setting, energy is no longer conserved along geodesics, rendering the second-order equations of motion irreducible to first-order forms. This constraint underscores the need to identify additional symmetries or devise innovative methods for calculating the black hole shadow.
Initial progress in this field included numerical approaches to shadow calculation, as demonstrated in early works Mishra et al. (2019); Koga et al. (2022). Analytically, significant strides were made in Vertogradov and Övgün (2024b), where a foundational relationship between energy conditions and the evolution of the black hole shadow was established. These pioneering efforts have laid the groundwork for understanding black hole shadows in the more complex setting of dynamical metrics, moving beyond the simpler static cases.
General relativistic corrections are crucial for the formation of both the photon sphere and the black hole shadow. These corrections determine the precise conditions under which the centrifugal force balances the effects of the gravitational field. In the following, we will demonstrate that the radius at which this balance occurs corresponds to the radius of the photon sphere, providing the foundation for its structure and observational properties.
This paper is structured as follows: Section 2 provides an in-depth discussion of geodesic motion in both static and dynamical spacetimes, establishing a connection between the energy of a particle and the generalized flux. Section 3 is dedicated to calculating the dynamical radius of the photon sphere and the black hole shadow, incorporating the effects of a changing spacetime geometry. Section 4 concludes the research, summarizing the key findings and their implications.
[…]
$\dot{M}$ $=$ $\frac{\partial M}{\partial v},$
$M^{\prime}$ $=$ $\frac{\partial M}{\partial r}.$
II Geodesic equations
We begin by exploring the foundational concepts of geodesic motion. Initially, we describe the motion of geodesics within a general spherically symmetric spacetime, emphasizing the fundamental principles governing such systems. Subsequently, we extend our analysis to the case of dynamical spacetimes, examining how the introduction of time dependence alters the behavior of geodesics.
By comparing the two models, we identify both their differences and similarities, shedding light on how dynamical effects influence geodesic motion.
II.1 Static case
The general line element in Eddington-Finkelstein coordinates $\{v,r,\theta,\varphi\}$ for a spherically symmetric spacetime can be expressed as:
$ds^{2}$ $=$ $-fdv^{2}+2\varepsilon dvdr+r^{2}d\Omega^{2},$
$f(r)$ $=$ $1-\frac{2M(r)}{r}.$ (1)
where:
  • $M(r)$ is the mass function that characterizes the gravitational field as a function of the radial coordinate r.
  • $\varepsilon=\pm 1$ is a parameter, often used to distinguish between different coordinate conventions, which has no impact on geodesic motion in the static case.
  • $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\varphi^{2}$ is the metric on the unit two-sphere, representing the angular part of the spacetime in spherical coordinates.
This form is particularly useful for studying geodesic motion because it separates the radial, temporal, and angular components of the spacetime, making it easier to analyze the dynamics in both static and dynamical cases a spacetime (II.1) is spherically-symmetric and static. The spacetime admits two Killing vectors, reflecting its symmetries under time translation and rotation. These symmetries result in two conserved quantities along geodesics: Energy per unit mass $(E)$, associated with the time-translation invariance. Angular momentum per unit mass $(L)$, stemming from rotational invariance.
$E$ $=$ $f\frac{dt}{d\lambda},$
$L$ $=$ $r^{2}\sin^{2}\theta\frac{d\varphi}{d\lambda},$ (2)
where the $\lambda$ is an affine parameter. In a spherically symmetric spacetime, the motion of particles and light can always be confined to a single plane due to the rotational symmetry. Without loss of generality, we choose the equatorial plane, where $\theta=\frac{\pi}{2}$. In this plane, the angular derivative terms vanish, i.e., $\frac{d^{2}\theta}{d\lambda^{2}}=\frac{d\theta}{d\lambda}=0$. This simplifies the analysis to motion in the equatorial plane. Using the geodesic condition $g_{ik}u^{i}u^{k}=\delta$, one finds
$\left(\frac{dr}{d\lambda}\right)^{2}$ $+$ $V_{eff}=0,$
$V_{eff}$ $=$ $f\left(\frac{L^{2}}{r^{2}}-\delta\right)-E^{2}.$ (3)
The parameter $\delta=0,\pm 1$ in the geodesic equation identifies the type of trajectory being considered: null geodesics $(\delta=0)$ describe the motion of light, timelike geodesics $(\delta=-1)$ correspond to the motion of massive particles, and spacelike geodesics $(\delta=+1)$ represent less common, hypothetical paths. The effective potential, $V_{\text{eff}}$, governs the motion and shapes the trajectory of the particle or photon, encapsulating the combined influence of gravitational and centrifugal forces. For circular orbits, where the radial coordinate r remains constant, two conditions must be met: the effective potential must vanish $(V_{\text{eff}}=0)$, and its first derivative with respect to r must also be zero $(\frac{dV_{\text{eff}}}{dr}=0)$. Additionally, the stability of the circular orbit is determined by the second derivative of the effective potential $(\frac{d^{2}V_{\text{eff}}}{dr^{2}})$. If $\frac{d^{2}V_{\text{eff}}}{dr^{2}}>0$, the orbit is stable, while $\frac{d^{2}V_{\text{eff}}}{dr^{2}}<0$ indicates an unstable orbit.
$\frac{d^{2}r}{d\lambda^{2}}=-\frac{1}{2}V^{\prime}_{eff},$ (4)
which leads to
$\frac{d^{2}r}{d\lambda^{2}}=\delta\frac{M-M^{\prime}r}{r^{2}}+\left(M^{\prime}r-3M\right)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}},$ (5)
where $M{\prime}$ is the derivative of the mass function with respect to $r$, and $M$ is the mass function itself. $L$ is the angular momentum per unit mass.
We identify three key terms that contribute to the effective forces governing motion in the spacetime:
  • Newtonian Attractive Force $(a_{n})$: This term represents the classical gravitational attraction, given by: $a_{n}$
    $a_{n}=\frac{M^{\prime}r-M}{r^{2}}.$ (6)
  • Centrifugal Repulsive Force $a_{c}$: This term accounts for the repulsive effect of angular momentum, expressed as:
    $a_{c}=\frac{L^{2}}{r^{3}}.$ (7)
  • General Relativistic Corrections $a_{g}$: This term arises from general relativistic effects and modifies the overall dynamics of the system. Its specific form will depend on the spacetime under consideration and encapsulates deviations from classical mechanics due to the curvature of spacetime.
    $a_{g}=\frac{M^{\prime}r-3M}{r^{4}}L^{2}.$ (8)
Hence, the geodesic equation (5), incorporating the identified force terms, can be expressed as:
$\frac{d^{2}r}{d\lambda^{2}}=-\delta a_{n}+a_{c}+a_{g}.$ (9)
It is important to note that the Newtonian attractive force, $a_{n}$, is present only for timelike or spacelike geodesic motions. The radius $r_{ph}$ is the radius of the photon sphere if the following conditions are satisfied:Claudel et al. (2001)
  • The effective radial acceleration is zero, indicating equilibrium:
    $V_{eff}\left(r_{ph}\right)=0,~{}~{}a_{g}\left(r_{ph}\right)+a_{c}\left(r_{ph}\right)=0.$ (10)
  • The second condition for the photon sphere radius, $r_{ph}$, follows from the requirement that the derivative of the effective potential vanishes at that radius, i.e.,$V^{\prime}_{eff}\left(r_{ph}\right)=0$. This ensures that the forces at $r_{ph}$ are in equilibrium.
Therefore, we can reformulate the definition of the photon sphere as follows:
A radius of a photon sphere is the radius at which the centrifugal repulsive force and general relativistic corrections are balanced.
This revised definition provides a clear physical interpretation and establishes a basis for extending the concept to a dynamical photon sphere, which will be discussed further below.
II.2 Dynamical case
We consider a spherically symmetric dynamical spacetime in Eddington-Finkelstein coordinates $v,r,\theta,\varphi$ described by the following line element:
$ds^{2}=-f(v,r)dv^{2}+2\varepsilon dvdr+r^{2}d\Omega^{2},$ (11)
and the function $f(v,r)$ represents the lapse function and is given by:
$f(v,r)=1-\frac{2M(v,r)}{r},$ (12)
with $M(v,r)$ as the mass function, which depends on both the Eddington time coordinate $v$ and the radial coordinate $r$. The parameter $\varepsilon=\pm 1$ distinguishes between ingoing $(\varepsilon=+1)$ and outgoing $(\varepsilon=-1)$ radiation. In a spherically symmetric spacetime, the symmetry allows us to simplify the analysis of geodesic motion by restricting it to the equatorial plane, where $\theta=\frac{\pi}{2}$.
To describe the geodesic equation, we start with the action $I$, which governs the motion of a particle or photon in a curved spacetime. The action is given by:
[…]
The resulting equations of motion for each coordinate are as follows:
$\frac{d\varphi}{d\lambda}=\frac{L}{r^{2}},$ (14)
$-\frac{1}{2}f^{\prime}\left(\frac{dv}{d\lambda}\right)^{2}+r\left(\frac{d\varphi}{d\lambda}\right)^{2}-\varepsilon\frac{d^{2}v}{d\lambda^{2}}=0,$ (15)
$\varepsilon\frac{d^{2}r}{d\lambda^{2}}=\frac{1}{2}\dot{f}\left(\frac{dv}{d\lambda}\right)^{2}+f\frac{d^{2}v}{d\lambda^{2}}+f^{\prime}\frac{dv}{d\lambda}\frac{dr}{d\lambda}.$ (16)
In the equations of motion derived from the action, $\lambda$ represents the affine parameter along the geodesic, which parameterizes the trajectory of a particle or photon. The conserved angular momentum per unit mass is denoted by $L$, a quantity arising from the rotational symmetry of the spacetime. A dot $\dot{}$ refers to the partial derivative with respect to the Eddington-Finkelstein time coordinate $v$. A dash ($\prime$) indicates the partial derivative with respect to the radial coordinate $r$.
Expressing $f\frac{d^{2}v}{d\lambda^{2}}$ from (15) by using (14), we obtain
$f\frac{d^{2}v}{d\lambda^{2}}=\varepsilon f\frac{L^{2}}{r^{3}}-\frac{1}{2}\varepsilon ff^{\prime}\left(\frac{dv}{d\lambda}\right)^{2}.$ (17)
The null geodesic condition $g_{ik}\frac{dx^{i}}{d\lambda}\frac{dx^{k}}{d\lambda}=0$ gives
$f^{\prime}\frac{dv}{d\lambda}\frac{dr}{d\lambda}=\frac{1}{2}\varepsilon ff^{\prime}-\frac{1}{2}\varepsilon f^{\prime}\frac{L^{2}}{r^{2}}\left(\frac{dv}{d\lambda}\right)^{2}.$ (18)
Now, substituting (17) and (18) into (16), we finaly obtain
$\frac{d^{2}r}{d\lambda^{2}}=\frac{1}{2}\varepsilon\dot{f}\left(\frac{dv}{d\lambda}\right)^{2}-\frac{1}{2}f^{\prime}\frac{L^{2}}{r^{2}}+f\frac{L^{2}}{r^{3}}.$ (19)
By substituting (12) we arrive at general form of geodesic equation
$\frac{d^{2}r}{d\lambda^{2}}=-\varepsilon\frac{\dot{M}}{r}\left(\frac{dv}{d\lambda}\right)^{2}+\left(M^{\prime}r-3M\right)\frac{L^{2}}{r^{4}}+\frac{L^{2}}{r^{3}}.$ (20)
The first term on the right hand-side of (20) represents a non-Newtonian gravitational force arising from the dynamics of the black hole, specifically its radiation or accretion processes, which alter the mass function $M(v,r)$ over time. This term captures the energy exchange between the black hole and its surroundings, distinguishing it from the static Newtonian force. The second term accounts for relativistic corrections predicted by general relativity, which modify the gravitational force significantly in strong-field regions, such as near the event horizon or the photon sphere. Lastly, the third term corresponds to the repulsive centrifugal force resulting from the angular momentum of the particle or photon, which opposes the inward pull of gravity and establishes equilibrium in circular motion Vertogradov and Kudryavcev (2023).
The generalized total apparent flux, $\Lambda$, associated with the black hole’s accretion rate, can be defined as Lindquist et al. (1965):
$\Lambda=\varepsilon\dot{M}\left(\frac{dv}{d\lambda}\right)^{2}\,,$ (21)
We refer to the first term on the right-hand side as the induced acceleration, denoted by $a_{i}\equiv-\frac{\Lambda}{r}$. Additionally, we denote the general relativistic correction and the repulsive centrifugal force as $a_{g}$ and $a_{c}$, respectively. With these definitions, Eq. (20) can be rewritten as:
$\frac{d^{2}r}{d\lambda^{2}}=a_{i}(v,r)+a_{g}(v,r)+a_{c}(v,r).$ (22)
In dynamical spacetimes, it is a well-known fact that the energy per unit mass of a particle is not conserved along null geodesics due to the time-dependent nature of the spacetime. However, one can still define this energy using the Lagrangian formalism.
$2\mathcal{L}=g_{ik}\frac{dx^{i}}{d\lambda}\frac{dx^{k}}{d\lambda},$ (23)
as follows
$E(v)=-\frac{\partial\mathcal{L}}{\partial\dot{v}}=f\frac{dv}{d\lambda}-\frac{dr}{d\lambda}.$ (24)
By using (14), (24) and null geodesic equation condition $g_{ik}u^{i}u^{k}=0$, one obtains
$\left(\frac{dr}{d\lambda}\right)^{2}+V_{eff}(v,r)=0,$
$V_{eff}(v,r)=f\frac{L^{2}}{r^{2}}-E^{2}(v).$ (25)
To express the energy as a function of the total generalized flux $\Lambda$, we start by differentiating the energy expression with respect to the affine parameter $\Lambda$. From the definition of energy (II.2):
$02\frac{dr}{d\lambda}\left(\frac{d^{2}r}{d\lambda^{2}}+\frac{1}{2}\frac{\partial V_{eff}}{\partial r}\right)+\frac{\partial V_{eff}}{\partial v}\frac{dv}{d\lambda}=$ 0 . (26)
However
$\frac{1}{2}\frac{\partial v_{eff}}{\partial r}=a_{g}-a_{c}.$ (27)
By using (20) we arrive at
$-2\frac{\Lambda}{r}\frac{dr}{d\lambda}+\frac{\partial V_{eff}}{\partial v}\frac{dv}{d\lambda}=0.$ (28)
Remembering the definition of (21) and using the fact that
$\frac{\partial V_{eff}}{\partial v}$ $=$ $-\frac{2\dot{M}L^{2}}{r^{3}}-2E\dot{E},$
$2\frac{dr}{d\lambda}\frac{dv}{d\lambda}$ $=$ $f\left(\frac{dv}{d\lambda}\right)^{2}-\frac{L^{2}}{r^{2}},$ (29)
we obtain
$\frac{d}{dv}\left(E^{2}\right)$ $=$ $-\frac{\Lambda}{r},$
[…]
Given that $\Lambda\propto\dot{M}$, the generalized flux is directly related to the rate of change of the black hole’s mass. When the spacetime transitions back to a static configuration, $\dot{M}\to 0$, implying that $\Lambda\to 0$. In this static limit, the dynamics simplify, and the energy per unit mass of the particle becomes conserved. This leads to: $\frac{dE}{d\lambda}\to 0\quad\Rightarrow\quad E^{2}=\text{constant}.$ In this scenario, the geodesic equation reduces to the form seen in a static spacetime, recovering the standard result for conserved energy, as in (II.2).
The total generalized flux $\Lambda$ can indeed be associated with the change in energy, offering insights into the relationship between energy dynamics and energy conditions. Specifically:
  • For a dissipative process, where energy decreases over time ($\dot{E}<0$), we have $\Lambda>0$, implying $\dot{M}>0$. This corresponds to an accretion process, where the black hole gains mass. In this case, the null energy condition (NEC) is satisfied.
  • Conversely, if the energy grows over time ($\dot{E}>0$), then $\Lambda<0$, which implies $\dot{M}<0$. This situation reflects a radiative process where the black hole loses mass. Here, the NEC is violated, as the energy flux is outward.
It is also worth noting that the second-order geodesic equations for timelike or spacelike trajectories differ from those for null trajectories by the inclusion of an additional term: the Newtonian attractive force, $a_{n}$.
The general form of the radial geodesic equation can be written as:
$\frac{d^{2}r}{d\lambda^{2}}=2E\dot{E}-\delta a_{n}+a_{g}+a_{c}.$ (31)
This equation unifies the contributions from energy changes, Newtonian gravity, relativistic corrections, and centrifugal forces, offering a comprehensive description of geodesic motion in both static and dynamical spacetimes.
III Dynamical black hole shadow
III.1 Dynamical photon sphere
To determine the radius of a dynamical photon sphere, we make the following two assumptions:
  1. 1.
    Slow Rate of Accretion: The generalized flux is small, $\Lambda\ll 1$, implying that the dynamical changes in the black hole’s mass are slow.
  2. 2.
    Radial Acceleration is Zero: The photon sphere is defined by equilibrium, i.e., $\frac{d^{2}r}{d\lambda^{2}}=0$.
To calculate the radius of the photon sphere, we introduce a book-keeping parameter $\alpha$ to facilitate an expansion. This parameter allows us to systematically account for small perturbations. We assume the radius of the photon sphere can be expressed as:
$r_{ph}=r_{0}+\alpha r_{1}(v),$ (32)
where:
  • $r_{0}$ is the radius of the photon sphere in the corresponding static spacetime.
  • $r_{1}(v)$ is a time-dependent perturbation that accounts for the dynamical effects introduced by the time dependence of the spacetime metric.
After performing the necessary expansions and calculations, the parameter $\alpha$ will be set to unity to recover the physical radius. This approach enables a perturbative treatment of the photon sphere’s radius in slowly evolving black hole spacetimes.
The geodesic equation (22) can be written as:
$\alpha a_{i}+a_{g}+a_{c}=0,$ (33)
and expanding with respect to $\alpha$, we find that $r_{0}$ satisfies the algebraic equation:
$a_{g}+a_{c}=0,$ (34)
and the first-order correction is given by:
$r_{1}=-\frac{a_{i}}{a^{\prime}_{g}+a^{\prime}_{c}}.$ (35)
In the paper Vertogradov and Övgün (2024b), we proved that if the weak energy condition is satisfied, then the radius of the photon sphere grows. Let us prove that $r_{1}<0$ only if the weak energy condition is violated. From the definition (21) and $a_{i}$, one can see that the weak energy condition demands $a_{i}\leq 0$ ($\dot{M}\geq 0$). Thus, we need to prove that the denominator of (35) is positive. From the definitions of $a_{g}$ and $a_{c}$, we find:
$a^{\prime}_{g}$ $=$ $\frac{L^{2}}{r^{5}}\left(M^{\prime\prime}r^{2}-6M^{\prime}r+12M\right),$
$a^{\prime}_{c}$ $=$ $-\frac{3L^{2}}{r^{4}}.$ (36)
Thus, one should demand the following inequality  [^1] Remember that we estimate (III.1) at $r=r_{0}$:
$M^{\prime\prime}r^{2}-6M^{\prime}r+12M-3r\geq 0.$ (37)
The dominant energy condition demands $-\frac{M^{\prime\prime}}{r}\leq 2\frac{M^{\prime}}{r^{2}}$. Applying this condition, we obtain:
$-8M^{\prime}r+12M-3r\geq 0$ (38)
The radius of a photon sphere should satisfy the following condition:
$6M-2M^{\prime}r-2r=0,$
$12M-3r-8M^{\prime}r=5r-4M\geq 0$ (39)
The last inequality holds because $r_{0}\geq\frac{3}{2}r_{h}$, where $r_{h}=2M(r_{h},v)$ is the radius of an apparent horizon. Thus, we prove that the denominator of (35) is always positive. This implies that if the weak energy condition is satisfied, then the radius of the photon sphere grows over time, which coincides with the earlier results obtained in Vertogradov and Övgün (2024a).
As an example, let us consider the simplest case of the Vaidya spacetime, which is given by:
$ds^{2}=-\left(1-\frac{2M(v)}{r}\right)dv^{2}+2dvdr+r^{2}d\Omega^{2}.$ (40)
The corresponding static spacetime is the Schwarzschild spacetime, for which the photon sphere radius is $r_{0}=3M_{0}$, where $M_{0}$ is the mass of the Schwarzschild black hole. In the case of accretion ($\dot{M}>0$), the induced acceleration satisfies $a_{i}<0$.
The derivatives of the relativistic correction $a_{g}$ and the centrifugal force $a_{c}$ are given by:
$a^{\prime}_{g}$ $=$ $\frac{12M(v)}{(3M_{0})^{5}}L^{2},$
$a^{\prime}_{c}$ $=$ $-\frac{3L^{2}}{(3M_{0})^{4}}.$ (41)
The first-order correction to the photon sphere radius, $r_{1}$, is then:
$r_{1}=-\frac{(3M_{0})^{5}a_{i}}{12M(v)L^{2}-9M_{0}L^{2}}\geq 0.$ (42)
One should note, that we consider an accretion process onto a black hole of the mass $M_{0}$. That’s why, $M(v)\geq M_{0}$ and the denominator is always positive. Thus, in the case of accretion, the radius of the photon sphere, given by:
$r_{ph}=r_{0}+r_{1}(v),$ (43)
is always increasing shown in Fig. 1. This result demonstrates that accretion leads to an outward shift of the photon sphere in the Vaidya spacetime.
Figure 1: The Figure shows the variation of $r_{ph}$ due to the parameter $M(v)$ with constant $L=1$ and $a_{i}=-0.1$. Hence $r_{ph}$ remains positive throughout, indicating a increase in the black hole’s photon sphere.
III.2 Dynamical shadow
In order to calculate the angular radius of a shadow seen by an observer, we use the tetrad field, which defines the local orthonormal basis for the observer. The components of the tetrad field are given by:
$e_{0}$ $=$ $\frac{1}{\sqrt{1-\frac{2M}{r}}}\frac{d}{dv},$
$e_{1}$ $=$ $e_{0}+\sqrt{1-\frac{2M}{r}}\frac{d}{dr},$
$e_{2}$ $=$ $\frac{1}{r}\frac{d}{d\theta},$
$e_{3}$ $=$ $\frac{1}{r\sin\theta}\frac{d}{d\varphi}.$ (44)
In the equatorial plane ($\theta=\frac{\pi}{2}$), the null geodesic is described by the coordinates $v(\lambda)$, $r(\lambda)$, and $\varphi(\lambda)$. The tangent vector to the null geodesic can be expressed as:
$\dot{v}\frac{d}{dv}+\dot{r}\frac{d}{dr}+\dot{\varphi}\frac{d}{d\varphi}=\xi\left(e_{0}+\cos\omega\,e_{1}-\sin\omega\,e_{3}\right).$ (45)
Here, $\xi$ is a scalar factor, and $\omega$ is the angle between the null geodesic and the radial direction in the observer’s rest frame. By comparing the coefficients of $\frac{d}{dr}$ and $\frac{d}{d\varphi}$, we can write:
$\dot{r}$ $=$ $\sqrt{1-\frac{2M}{r}}\cos\omega,$
$\dot{\varphi}$ $=$ $-\frac{1}{r}\sin\omega.$ (46)
To proceed further, we need an expression for $\frac{dr}{d\varphi}$. Using our ansatz (II.2), along with (III.2), and (14), we find:
$\left(\frac{dr}{d\varphi}\right)^{2}$ $=$ $r^{2}f\cot^{2}\omega,$
$\sin^{2}\omega_{sh}$ $=$ $\frac{L^{2}}{E^{2}(v)}\frac{f\left(r_{o}\right)}{r_{o}^{2}},$ (47)
where $\sin^{2}\omega_{sh}$ represents the angular size of the shadow as seen by the observer located at $r_{o}$.
Hence, the angular size of the shadow decreases if $\dot{E}>0$ and increases if $\dot{E}<0$. One can assert the following: In a dynamical black hole, the shadow grows if light loses energy while propagating near the black hole and decreases if light gains energy while propagating near the black hole. We know, that an observer obtains light if the following condition is held
$\frac{L^{2}}{E^{2}(v)}\geq\frac{r^{2}_{ph}}{f(r_{ph})}.$ (48)
Hence, we can estimate the angular size of a shadow as
$\sin^{2}\omega_{sh}=\frac{r^{2}_{ph}f(r_{o})}{r_{o}^{2}f(r_{ph})}.$ (49)
Also one can write
$E^{2}(v)=E^{2}_{0}+\mu(v).$ (50)
Here $E_{0}$ is initial energy of a particle and $\mu(v)$ is an arbitrary function of time. $\mu(v)>0$ if null energy condition is violated and $\mu(v)<0$ if it is held. Denoting $\gamma(v)\equiv\frac{\mu(v)}{E^{2}_{0}}$, one can write
$\sin^{2}\omega_{sh}=\frac{b_{0}^{2}}{1+\gamma(v)}\frac{f(r_{0})}{r_{0}^{2}}.$ (51)
Here $b_{0}$ is a shadow radius of corresponding static spacetime.
For $r_{o}\gg M$, the term $\frac{2M}{r_{o}}$ becomes negligible compared to $1$. Hence, the metric function at the observer radius simplifies as: $f(r_{o})=1-\frac{2M}{r_{o}}\approx 1$. Substituting this approximation into the expression for $\sin^{2}\omega_{sh}$ gives approximately:
$\omega_{sh}\approx\frac{3}{r_{o}}\sqrt{\frac{3M}{1-\nu v}\cdot}.$ (52)
For Vaidya spacetime one can take $b_{0}^{2}=27M$ and $\gamma(v)=-\nu v,~{}~{}\nu>0$.
The plot in 2 shows that as $v$ increases, $\omega_{sh}$ grows due to the factor $(1-\nu v)^{-1/2}$, reflecting the time-dependence introduced by the Vaidya spacetime.
Figure 2: The Figure shows the plot of $\omega_{sh}(v)$ as a function of the retarded time $v$ for the Vaidya spacetime, using the approximation for $r_{o}\gg M$. The parameters are $r_{o}=10$, $M=1$, and $\nu=0.1$.
Figure 3: The Figure shows the visualization of the black hole shadow for different combinations of mass $(M=10,20)$ and retarded time $(v=2,5)$. Each subplot corresponds to a specific pair of $M$ and $v$, showing how the shadow radius change with these parameters.
The figure 3 illustrates the black hole shadow for varying combinations of mass $(M)$ and retarded time $(v)$, showcasing how these parameters influence the shadow’s size. As $M$ increases, the shadow radius grows significantly, consistent with the theoretical expectation that the shadow size scales with the black hole’s mass due to stronger gravitational lensing. For a fixed $M$, an increase in $v$ slightly enlarges the shadow radius, reflecting the dynamic nature of the Vaidya spacetime, where the metric evolves with time $v$. The mass has a dominant effect on the shadow size, while the time evolution modulates it further, demonstrating the interplay between static and dynamic factors in determining the observable shadow characteristics. This highlights the potential to use shadow observations to probe both the mass and dynamical properties of black holes.
IV Conclusion
A real astrophysical black hole is surrounded by an accretion disc. As a result of constant accretion onto the black hole, its mass changes over time. Consequently, the spacetime around the black hole becomes time-dependent. This time-dependence should, in turn, influence observable manifestations of the black hole, such as its shadow. It has previously been shown that the evolution of the shadow is intricately connected to the energy conditions of the spacetime Vertogradov and Övgün (2024b). Variations in the mass of the black hole, driven by accretion or radiation, directly affect the size and shape of the shadow, making it a potential tool to study dynamic processes in astrophysical black holes.
Recently, it has been shown that the shadow also depends on the structure of the apparent horizon Vertogradov (2024a). This connection implies that by studying the dynamical shadow, we can gain significant insights into the processes occurring near the horizon of a black hole. However, the analysis of dynamical processes is inherently challenging due to the absence of a conserved energy for the particle in a time-dependent spacetime. This lack of conserved energy complicates the equations of motion, preventing the reduction of the second-order equations to first-order ones, thereby making the study of dynamical black holes mathematically more complex.
In our work, we analyzed the second-order geodesic equation and expressed the radial component in terms of fictitious forces. In a static metric, the radius of the photon sphere is determined by the balance between the centrifugal force and the general relativistic corrections. By extending this framework to a dynamical spacetime, we assumed that the radius of the dynamical photon sphere is established by the balance of three forces: the induced force (arising from the black hole’s mass variation), the centrifugal force, and the general relativistic corrections. Using this assumption, we derived an approximate expression for the evolution of the photon sphere radius, capturing its dependency on the dynamic properties of the black hole and the surrounding spacetime.
Then, we examined the formation of the shadow of a dynamical black hole, expressing the change in the energy of light through the generalized flux. We demonstrated that the shadow of the black hole increases if the light loses energy and decreases if the light gains energy during its propagation near the black hole.
An intriguing observation is that when the radius of the photon sphere coincides with the radius of the NEC (Null Energy Condition) horizon Vertogradov (2024a), the dynamical shadow does not evolve, even though both the photon sphere and the NEC horizon radii are changing. This result highlights a surprising decoupling between the evolution of these radii and the observable shadow in this specific configuration.
We have developed a method for calculating the dynamical shadow of a black hole under the assumption of a small accretion rate. From an astrophysical perspective, the changes caused by accreting matter typically do not significantly affect the mass of the black hole. Consequently, the evolution of the photon sphere radius and the shadow of the black hole is expected to be slow and largely undetectable by modern experimental tools.
The proposed method represents a foundational step toward the analytical calculation of the shadow of a dynamical black hole in the general case. Its refinement will enable the description of the shadow in more extreme scenarios, such as the gravitational collapse of a star—a process so rapid that the evolution and formation of the shadow could become observable. This capability would provide profound insights into the dynamic processes involved in black hole formation and their observational signatures.
Acknowledgements.
V. Vertogradov thanks the Basis Foundation (grant number 23-1-3-33-1) for the financial support. A. Ö. would like to acknowledge the contribution of the COST Action CA21106 - COSMIC WISPers in the Dark Universe: Theory, astrophysics and experiments (CosmicWISPers) and the COST Action CA22113 - Fundamental challenges in theoretical physics (THEORY-CHALLENGES). We also thank TUBITAK and SCOAP3 for their support.